# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 3.2 CBPs with a known control function
If $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi^{(k)}$ have the same offspring mean $m$ and variance $\sigma^{2}$, then
$||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-1/2}).$
@@
We can note the resemblance of Theorem 3.4 to [23, Theorem 2] on the non-existence of consistent estimators for BGWPs (see also Theorem 2.1). Similarly, Theorem 3.4 does not tell us whether we can estimate $m$ and $\sigma^{2}$—only that we cannot consistently estimate anything other than functions of $m$ and $\sigma^{2}$.
We now show that it is possible to construct consistent estimators for $m$ and $\sigma^{2}$ for processes in $\Pi^{(k)}$. There are two cases of interest: first, when $m$ is known and $\sigma^{2}$ unknown, and second, when both $m$ and $\sigma^{2}$ are unknown. In the first case, we propose
$\bar{\sigma}_{n}^{2}:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-m\cdot\varepsilon(Z_{k-1}))^{2}-m^{2}\cdot\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}$
@@
$\hat{m}_{n}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{Z_{k}}{\varepsilon(Z_{k-1})},$
$\hat{\sigma}_{n}^{2}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-\tilde{m}_{n}\cdot\varepsilon(Z_{k-1}))^{2}-\tilde{m}_{n}^{2}\cdot\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}$
@@
Under mild assumptions, we can show that $\hat{m}_{n}$, $\bar{\sigma}_{n}^{2}$, and $\hat{\sigma}_{n}^{2}$ are all consistent estimators.
@@
Let $\{Z_{n},z_{0}\}\in\Pi^{(k)}$, and assume $\varepsilon(z)>0$ for all $z\in\mathbbm{N}_{1}$. Then, on the set $\{Z_{n}\to\infty\}$,
@@
If we further assume that $\sup_{z\geq 1}\left\{\frac{|\upsilon(z)|}{z}\right\}\leq c$ and $\sup_{z\geq 1}\Big{\{}\frac{\mathbbm{E}(\phi(z)-\varepsilon(z))^{4}}{z^{2}}\Big{\}}\leq d$ for positive constants $c$ and $d$, and that $\mathbbm{E}(\xi-m)^{4}$ is finite, then
@@ 3.3 CBPs with an unknown control function
In most cases, we would not expect that the distribution of the control function is known beforehand; hence, it needs to be estimated alongside the offspring distribution. Recall that a control function $\{\phi(z)\}_{z\in\mathbbm{N}_{0}}$ is specified by a countably infinite set of random variables, one for each population size $z$. To make the problem of estimating the distribution of $\phi(\cdot)$ tractable, we require some regularity conditions. First, to ensure that the distribution of $\phi(z)$ converges to the normal distribution as $z\to\infty$, we assume that the control function is linearly-divisible:
@@
The control function $\{\phi(z)\}_{z\in\mathbbm{N}_{0}}$ is linearly-divisible if there exists a function $l:\mathbbm{N}_{0}\to\mathbbm{N}_{0}$ such that $l(z)=\Theta(z)$ (i.e., $0<\liminf_{z\to\infty}l(z)/z$ and $\limsup_{z\to\infty}l(z)/z<\infty$) and, for each $z\in\mathbbm{N}_{0}$, there exist a set of i.i.d. random variables $\{\chi_{i}^{(z)}\}_{1\leq i\leq l(z)}$ such that $\phi(z)\stackrel{{\scriptstyle d}}{{=}}\sum_{i=1}^{l(z)}\chi_{i}^{(z)}$.
Recall that (6) provides a sufficient condition for a supercritical CBP to have a positive probability of unbounded growth. Here, we need to strengthen this condition to include a bound on the growth rate of the third central moment of the control function, $\upsilon(z):=\mathbbm{E}[\phi(z)-\mathbbm{E}\phi(z)]^{3}$. In particular, we assume that there exist constants $a,b,c\geq 0$ such that
$\sup_{z\geq 1}\left\{\frac{\varepsilon(z)}{z}\right\}\leq a,\quad\sup_{z\geq 1}\left\{\frac{\nu^{2}(z)}{z}\right\}\leq b,\quad\text{and}\quad\sup_{z\geq 1}\left\{\frac{|\upsilon(z)|}{z}\right\}\leq c.$ (8)
@@
for some $r<1$, then there exists $q>0$ such that
$||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q}).$
We also obtain the analogue of Theorem 3.4:
@@
$(X_{1}|X_{0}=z)$ $\stackrel{{\scriptstyle d}}{{\approx}}\mathbbm{E}(X_{1}|X_{0}=z)+\sqrt{\text{Var}(X_{1}|X_{0}=z)}\,\mathcal{N}(0,1)$ (10)
$(Z_{1}|Z_{0}=z)$ $\stackrel{{\scriptstyle d}}{{\approx}}\mathbbm{E}(Z_{1}|Z_{0}=z)+\sqrt{\text{Var}(Z_{1}|Z_{0}=z)}\,\mathcal{N}(0,1).$
@@
$(Z_{1}|Z_{0}=z)\stackrel{{\scriptstyle d}}{{\approx}}$
$\quad\mathbbm{E}(X_{1}|X_{0}=z)+\sqrt{\text{Var}(X_{1}|X_{0}=z)}\,\mathcal{N}\left(\frac{\mathbbm{E}(Z_{1}|Z_{0}=z)-\mathbbm{E}(X_{1}|X_{0}=z)}{\sqrt{\text{Var}(X_{1}|X_{0}=z)}}\,,\,\frac{\text{Var}(Z_{1}|Z_{0}=z)}{\text{Var}(X_{1}|X_{0}=z)}\right).$
@@
On the other hand, if we would like to estimate $\boldsymbol{\theta}^{(2)}=(g,\,h):=(m\alpha,\,\sigma^{2}\alpha+m^{2}\beta)$ (where $g$ is the mean growth rate of the process), then (9) does not hold for any $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi^{(u*)}$ with $\boldsymbol{\theta}^{(2)}_{Z}\neq\boldsymbol{\theta}^{(2)}_{X}$. We let
$\hat{g}_{n}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{Z_{k}}{Z_{k-1}},$
$\bar{h}_{n}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-m\alpha\cdot Z_{k-1})^{2}}{Z_{k-1}},\text{ and}$
$\hat{h}_{n}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-\tilde{g}_{n}\cdot Z_{k-1})^{2}}{Z_{k-1}},$
@@ 3.4 CBPs with observed progenitor numbers
$||\mathcal{L}_{(\phi_{Z}(Z_{0}),\,Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi_{X}(X_{0}),\,X_{1})|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q}).$
@@
Consider again the class of supercritical CBPs $\{Z_{n},z_{0}\}\in\Pi^{(u*)}$, with control functions $\phi(\cdot)$ satisfying $\varepsilon(z)=\alpha z$ and $\nu^{2}(z)=\beta z$ for $\alpha,\beta>0$. Recall from the previous section that if only the population sizes are observed at each generation, then $\boldsymbol{\theta}^{(1)}:=(m,\,\sigma^{2},\,\alpha,\,\beta)$ cannot be estimated consistently. Under the current observation scheme, we consider the estimators
$\hat{m}_{n}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{Z_{k}}{\phi(Z_{k-1})},$
$\bar{\sigma}_{n}^{2}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-m\cdot\phi(Z_{k-1}))^{2}}{\phi(Z_{k-1})},$
$\hat{\sigma}_{n}^{2}$ $:=\frac{1}{n}\sum_{k=1}^{n}\frac{(Z_{k}-\tilde{m}_{n}\cdot\phi(Z_{k-1}))^{2}}{\phi(Z_{k-1})},$
$\hat{\alpha}_{n}$ $:=\frac{1}{n}\sum_{k=0}^{n-1}\frac{\phi(Z_{k})}{Z_{k}},$
$\bar{\beta}_{n}$ $:=\frac{1}{n}\sum_{k=0}^{n-1}\frac{(\phi(Z_{k})-\alpha\cdot Z_{k})^{2}}{Z_{k}},\quad\text{and}$
$\hat{\beta}_{n}$ $:=\frac{1}{n}\sum_{k=0}^{n-1}\frac{(\phi(Z_{k})-\tilde{\alpha}_{n-1}\cdot Z_{k})^{2}}{Z_{k}},$
@@ 4 Concluding remarks
As mentioned in Section 1, a common rule of thumb in population ecology is that demographic and environmental stochasticity should not be simultaneously estimated from a single trajectory of population size counts. This rule is supported by Theorems 3.8 and its application to the processes in $\Pi^{(u*)}$ with parameter $\boldsymbol{\theta}^{(1)}:=(m,\,\sigma^{2},\,\alpha,\,\beta)$. Our results suggest three different ways to address this limitation: (i) Use an independent data source, or expert knowledge, to estimate environmental stochasticity (i.e., the control function) first, then estimate demographic stochasticity (i.e., $m$ and $\sigma^{2}$) while treating the distribution of the control function as known (supported by Theorem 3.5 in setting S1); (ii) Rely on expert knowledge on the species to estimate the demographic parameters, and use population size counts to estimate the control function parameters only (for example $\alpha$ and $\beta$, as supported by Theorem 3.10 and commentary below); (iii) Collect additional data beyond a single trajectory of population size counts, as in setting S3 (Theorem 3.13).
@@ 5 Proofs
Let $\{Z_{n}\}_{n\in\mathbbm{N}_{0}}$ and $\{X_{n}\}_{n\in\mathbbm{N}_{0}}$ be two Markov chains taking values on $\mathbbm{N}_{0}$. Then, for $j,k\in\mathbbm{N}_{1}$ such that $j<k$, if there exists a monotonically decreasing function $\boldsymbol{\mathfrak{K}}:[M,\infty)\to\mathbbm{R}_{\geq 0}$, $M\in\mathbbm{N}_{0}$, such that
$||\mathcal{L}_{(Z_{j+1},\dots,Z_{k})|Z_{j}=u_{j}}-\mathcal{L}_{(X_{j+1},\dots,X_{k})|X_{j}=u_{j}}||_{TV}\leq\boldsymbol{\mathfrak{K}}(u_{j})\quad\forall u_{j}\geq M,$ (15)
we have that
$||\mathcal{L}_{(Z_{j},\dots,Z_{k})|Z_{j-1}=u_{j-1}}-\mathcal{L}_{(X_{j},\dots,X_{k})|X_{j-1}=u_{j-1}}||_{TV}$
$\leq||\mathcal{L}_{Z_{j}|Z_{j-1}=u_{j-1}}-\mathcal{L}_{X_{j}|X_{j-1}=u_{j-1}}||_{TV}+\mathbbm{P}(Z_{j}\leq N|Z_{j-1}=u_{j-1})+\boldsymbol{\mathfrak{K}}(N+1)$
@@
Proof. We write $p_{Z_{n}}(i,j):=\mathbbm{P}(Z_{n}=j|Z_{n-1}=i)$ and $p_{X_{n}}(i,j):=\mathbbm{P}(X_{n}=j|X_{n-1}=i)$. Then, from the sum representation of the total variation distance [22, Proposition 4.2] and the triangle inequality, we see that
$||\mathcal{L}_{(Z_{j},\dots,Z_{k})|Z_{j-1}=u_{j-1}}-\mathcal{L}_{(X_{j},\dots,X_{k})|X_{j-1}=u_{j-1}}||_{TV}$
@@
$=||\mathcal{L}_{Z_{j}|Z_{j-1}=u_{j-1}}-\mathcal{L}_{X_{j}|X_{j-1}=u_{j-1}}||_{TV}$
$\quad+\sum_{u_{j}=0}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\cdot||\mathcal{L}_{(Z_{j+1},\dots,Z_{k})|Z_{j}=u_{j}}-\mathcal{L}_{(X_{j+1},\dots,X_{k})|X_{j}=u_{j}}||_{TV}$
$\leq||\mathcal{L}_{Z_{j}|Z_{j-1}=u_{j-1}}-\mathcal{L}_{X_{j}|X_{j-1}=u_{j-1}}||_{TV}$
$\quad+\sum_{u_{j}=0}^{N}p_{Z_{j}}(u_{j-1},u_{j})\cdot 1+\sum_{u_{j}=N+1}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\cdot\boldsymbol{\mathfrak{K}}(u_{j})$
$\leq||\mathcal{L}_{Z_{j}|Z_{j-1}=u_{j-1}}-\mathcal{L}_{X_{j}|X_{j-1}=u_{j-1}}||_{TV}$
$\quad+\mathbbm{P}(Z_{j}\leq N|Z_{j-1}=u_{j-1})+\sum_{u_{j}=N+1}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\cdot\boldsymbol{\mathfrak{K}}(u_{j}).$
@@
$\sum_{u_{j}=N+1}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\cdot\boldsymbol{\mathfrak{K}}(u_{j})\leq\boldsymbol{\mathfrak{K}}(N+1)\cdot\sum_{u_{j}=N+1}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\leq\boldsymbol{\mathfrak{K}}(N+1).\$
Combining these two inequalities then yields our desired result. $\blacksquare$
@@
for constants $n\in\mathbbm{N}_{1}$, $c_{1},\dots,c_{n}>0$, $q_{1},\dots,q_{n}>0$, and $b>1$. Then
$\lim_{z\to\infty}\lim_{k\to\infty}f_{k}(z)=0.$
Proof. For a given $k\in\mathbbm{N}_{1}$, we can expand $f_{k}$ iteratively to see that
$f_{k}(z)$ $=\sum_{i=1}^{n}\frac{c_{i}}{z^{q_{i}}}\cdot\sum_{j=0}^{k-2}\big{(}b^{-q_{i}}\big{)}^{j}+f_{1}(b^{k-1}\cdot z)$
$=\sum_{i=1}^{n}\frac{c_{i}}{z^{q_{i}}}\cdot\frac{1-b^{-(k-1)q_{i}}}{1-b^{-q_{i}}}+f_{1}(b^{k-1}\cdot z).$
Since $b>1$ and $q_{i}>0$ for all $i\in[n]$, each $b^{-q_{i}}<1$. Therefore, given that $\lim_{x\to\infty}f_{1}(x)=0$, we have that
@@
from which our desired result follows by letting $z\to\infty$. $\blacksquare$
@@
Proof of Lemma 3.1. Since $\{Z_{n},z_{0}\}$ is assumed to be supercritical, $\liminf_{z\to\infty}\tau_{Z}(z)>1$. Hence for any $t$ such that $1<t<\liminf_{z\to\infty}\tau_{Z}(z)$, there exists $M_{1}\in\mathbbm{N}_{1}$ such that for all $z\geq M_{1}$, $\varepsilon(z)\cdot m>t\cdot z$. In addition, if $||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})$, then there exists $s>0$ such that for all $z\geq M_{2}$, $M_{2}\in\mathbbm{N}_{1}$, $||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}\leq s\cdot z_{0}^{-q}$.
@@
$||\mathcal{L}_{(Z_{j},\dots,Z_{k})|Z_{j-1}=u_{j-1}}-\mathcal{L}_{(X_{j},\dots,X_{k})|X_{j-1}=u_{j-1}}||_{TV}\leq\boldsymbol{\mathfrak{K}}_{k-j+1}(u_{j-1}),$ (16)
for a decreasing function $\boldsymbol{\mathfrak{K}}_{k-j+1}:[M_{3},\infty)\to\mathbbm{R}_{\geq 0}$ given by $\boldsymbol{\mathfrak{K}}_{1}(z):=s\cdot z^{-q}$ and for $j<k$, $\boldsymbol{\mathfrak{K}}_{k-j+1}(z):=s\cdot z^{-q}+\frac{(a\sigma^{2}+bm^{2})}{(1-\alpha)^{2}t^{2}\cdot z}+\boldsymbol{\mathfrak{K}}_{k-j}(\alpha t\cdot z)$, where $\alpha\in(1/t,1)$.
@@
$||\mathcal{L}_{Z_{k}|Z_{k-1}=u_{k-1}}-\mathcal{L}_{X_{k}|X_{k-1}=u_{k-1}}||_{TV}\leq\frac{s}{u_{k-1}^{q}}$
for any $u_{k-1}\geq M_{3}$.

### v2

@@ 3.2 CBPs with a known control function
If $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi^{(k)}$ have the same offspring mean $m$ and variance $\sigma^{2}$, then $\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-1/2})$.
@@
We can note the resemblance of Theorem 3.4 to [25, Theorem 2] on the non-existence of consistent estimators for BGWPs (see also Theorem 2.1). Similarly, Theorem 3.4 does not tell us whether we can estimate $m$ and $\sigma^{2}$—only that we cannot consistently estimate anything other than functions of $m$ and $\sigma^{2}$.
We now show that it is possible to construct consistent estimators for $m$ and $\sigma^{2}$ for processes in $\Pi^{(k)}$. There are two cases of interest: first, when $m$ is known and $\sigma^{2}$ unknown, and second, when both $m$ and $\sigma^{2}$ are unknown. Let $I^{\varepsilon}_{n}:=\{k\in\{1,\dots,n\}:\varepsilon(Z_{k-1})>0\}$. In the first case, we propose
$\bar{\sigma}_{n}^{2}:=\frac{1}{|I^{\varepsilon}_{n}|}\sum_{k\in I^{\varepsilon}_{n}}\frac{(Z_{k}-m\cdot\varepsilon(Z_{k-1}))^{2}-m^{2}\cdot\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}$
@@
$\hat{m}_{n}$ $:=\frac{1}{|I^{\varepsilon}_{n}|}\sum_{k\in I^{\varepsilon}_{n}}\frac{Z_{k}}{\varepsilon(Z_{k-1})},$ $\quad\hat{\sigma}_{n}^{2}$ $:=\frac{1}{|I^{\varepsilon}_{n}|}\sum_{k\in I^{\varepsilon}_{n}}\frac{(Z_{k}-\tilde{m}_{n}\cdot\varepsilon(Z_{k-1}))^{2}-\tilde{m}_{n}^{2}\cdot\nu^{2}(Z_{k-1})}{\varepsilon(Z_{k-1})}$
@@
Under mild assumptions, we can show that $\hat{m}_{n}$, $\bar{\sigma}_{n}^{2}$, and $\hat{\sigma}_{n}^{2}$ are all consistent estimators. In the next theorem and throughout the rest of the paper, we denote the third central moment of the control function by $\iota(z):=\mathbbm{E}(\phi(z)-\mathbbm{E}\phi(z))^{3}$.
@@
Let $\{Z_{n},z_{0}\}\in\Pi^{(k)}$. Then, on the set $\{Z_{n}\to\infty\}$,
@@
If we further assume that $\sup_{z\geq 1}\left\{\frac{|\iota(z)|}{z}\right\}\leq c$ where and $\sup_{z\geq 1}\Big{\{}\frac{\mathbbm{E}(\phi(z)-\varepsilon(z))^{4}}{z^{2}}\Big{\}}\leq d$ for positive constants $c$ and $d$, and that $\mathbbm{E}(\xi-m)^{4}$ is finite, then
@@ 3.3 CBPs with an unknown control function
In most cases, we would not expect that the distribution of the control function is known beforehand; hence, it needs to be estimated alongside the offspring distribution. Recall that a control function $\{\phi(z)\}_{z\in\mathbbm{N}_{0}}$ is specified by a countably infinite set of random variables, one for each population size $z$. To make the problem of estimating the distribution of $\phi(\cdot)$ tractable, we require some regularity conditions. First, to ensure that the (appropriately scaled) distribution of $\phi(z)$ converges to the normal distribution as $z\to\infty$, we assume that the control function is linearly-divisible:
@@
The control function $\{\phi(z)\}_{z\in\mathbbm{N}_{0}}$ is linearly-divisible if there exists a function $l:\mathbbm{N}_{0}\to\mathbbm{N}_{0}$ such that $l(z)=\Theta(z)$ (i.e. $0<\liminf_{z\to\infty}l(z)/z$ and $\limsup_{z\to\infty}l(z)/z<\infty$) and, for each $z\in\mathbbm{N}_{0}$, there exist a set of i.i.d. random variables $\{\chi_{i}^{(z)}\}_{1\leq i\leq l(z)}$ such that $\phi(z)\stackrel{{\scriptstyle d}}{{=}}\sum_{i=1}^{l(z)}\chi_{i}^{(z)}$.
Recall that (6) provides a sufficient condition for a supercritical CBP to have a positive probability of unbounded growth. Here, we need to strengthen this condition to include a bound on the growth rate of the third central moment of the control function. In particular, we assume that there exist constants $a,b,c\geq 0$ such that
$\sup_{z\geq 1}\left\{\frac{\varepsilon(z)}{z}\right\}\leq a,\quad\sup_{z\geq 1}\left\{\frac{\nu^{2}(z)}{z}\right\}\leq b,\quad\text{and}\quad\sup_{z\geq 1}\left\{\frac{|\iota(z)|}{z}\right\}\leq c.$ (8)
@@
for some $r<1$, then there exists $q>0$ such that $\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})$.
By feeding Lemma 3.7 into Proposition 3.2, we obtain the analogue of Theorem 3.4:
@@
$(X_{1}|X_{0}=z)$ $\stackrel{{\scriptstyle d}}{{\approx}}\mathbbm{E}(X_{1}|X_{0}=z)+\sqrt{\text{Var}(X_{1}|X_{0}=z)}\cdot\text{N}(0,1)$ (10)
$(Z_{1}|Z_{0}=z)$ $\stackrel{{\scriptstyle d}}{{\approx}}\mathbbm{E}(Z_{1}|Z_{0}=z)+\sqrt{\text{Var}(Z_{1}|Z_{0}=z)}\cdot\text{N}(0,1).$
@@
$(Z_{1}|Z_{0}=z)\stackrel{{\scriptstyle d}}{{\approx}}\mathbbm{E}(X_{1}|X_{0}=z)+\sqrt{\text{Var}(X_{1}|X_{0}=z)}\cdot N\left(\frac{\mathbbm{E}(Z_{1}|Z_{0}=z)-\mathbbm{E}(X_{1}|X_{0}=z)}{\sqrt{\text{Var}(X_{1}|X_{0}=z)}}\,,\,\frac{\text{Var}(Z_{1}|Z_{0}=z)}{\text{Var}(X_{1}|X_{0}=z)}\right).$
@@
On the other hand, if we would like to estimate $\boldsymbol{\theta}^{(2)}=(g,\,h):=(m\alpha,\,\sigma^{2}\alpha+m^{2}\beta)$ (where $g$ is the mean growth rate of the process), then (9) does not hold for any $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi^{(u*)}$ with $\boldsymbol{\theta}^{(2)}_{Z}\neq\boldsymbol{\theta}^{(2)}_{X}$. With $I_{n}:=\{k\in\{1,\dots,n\}:Z_{k-1}>0\}$, we let
$\hat{g}_{n}:=\frac{1}{|I_{n}|}\sum_{k\in I_{n}}\frac{Z_{k}}{Z_{k-1}},$
$\bar{h}_{n}:=\frac{1}{|I_{n}|}\sum_{k\in I_{n}}\frac{(Z_{k}-m\alpha\cdot Z_{k-1})^{2}}{Z_{k-1}},\quad\text{ and}\quad\hat{h}_{n}:=\frac{1}{|I_{n}|}\sum_{k\in I_{n}}\frac{(Z_{k}-\tilde{g}_{n}\cdot Z_{k-1})^{2}}{Z_{k-1}},$
@@ 3.4 CBPs with observed progenitor numbers
$\mathcal{L}_{(\phi_{Z}(Z_{0}),\,Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi_{X}(X_{0}),\,X_{1})|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q}).$ (15)
With our new observation scheme, Lemma 3.1 and Proposition 3.2 cannot be directly applied. However, close equivalents of these results exist, and together with Lemma 3.11 lead to the following result:
@@
Consider again the class of supercritical CBPs $\{Z_{n},z_{0}\}\in\Pi^{(u*)}$, with control functions $\phi(\cdot)$ satisfying $\varepsilon(z)=\alpha z$ and $\nu^{2}(z)=\beta z$ for $\alpha,\beta>0$. Recall from the previous section that if only the population sizes are observed at each generation, then $\boldsymbol{\theta}^{(1)}:=(m,\,\sigma^{2},\,\alpha,\,\beta)$ cannot be estimated consistently. Under the current observation scheme and with $I^{\phi}_{n}:=\{k\in\{1,\dots,n\}:\phi(Z_{k-1})>0\}$ and $I^{-}_{n}:=\{k\in\{0,\dots,n-1\}:Z_{k}>0\}$, we consider the estimators
$\hat{m}_{n}$ $:=\frac{1}{|I^{\phi}_{n}|}\sum_{k\in I^{\phi}_{n}}\frac{Z_{k}}{\phi(Z_{k-1})},$ $\quad\hat{\alpha}_{n}$ $:=\frac{1}{|I^{-}_{n}|}\sum_{k\in I^{-}_{n}}\frac{\phi(Z_{k})}{Z_{k}},$
$\bar{\sigma}_{n}^{2}$ $:=\frac{1}{|I^{\phi}_{n}|}\sum_{k\in I^{\phi}_{n}}\frac{(Z_{k}-m\cdot\phi(Z_{k-1}))^{2}}{\phi(Z_{k-1})},$ $\quad\bar{\beta}_{n}$ $:=\frac{1}{|I^{-}_{n}|}\sum_{k\in I^{-}_{n}}\frac{(\phi(Z_{k})-\alpha\cdot Z_{k})^{2}}{Z_{k}},$
$\hat{\sigma}_{n}^{2}$ $:=\frac{1}{|I^{\phi}_{n}|}\sum_{k\in I^{\phi}_{n}}\frac{(Z_{k}-\tilde{m}_{n}\cdot\phi(Z_{k-1}))^{2}}{\phi(Z_{k-1})},$ $\quad\hat{\beta}_{n}$ $:=\frac{1}{|I^{-}_{n}|}\sum_{k\in I^{-}_{n}}\frac{(\phi(Z_{k})-\tilde{\alpha}_{n-1}\cdot Z_{k})^{2}}{Z_{k}},$
@@ 4 Concluding remarks
As mentioned in Section 1, a common rule of thumb in population ecology is that demographic and environmental stochasticity should not be simultaneously estimated from a single trajectory of population size counts. This rule is supported by Theorem 3.8 and its application to the processes in $\Pi^{(u*)}$ with parameter $\boldsymbol{\theta}^{(1)}:=(m,\,\sigma^{2},\,\alpha,\,\beta)$. Our results suggest three different ways to address this limitation: (i) Use an independent data source, or expert knowledge, to estimate environmental stochasticity (i.e., the control function) first, then estimate demographic stochasticity (i.e., $m$ and $\sigma^{2}$) while treating the distribution of the control function as known (supported by Theorem 3.5 in setting S1); (ii) Rely on expert knowledge on the species to estimate the demographic parameters, and use population size counts to estimate the control function parameters only (for example $\alpha$ and $\beta$, as supported by Theorem 3.10 and commentary below); (iii) Collect additional data beyond a single trajectory of population size counts, as in setting S3 (Theorem 3.13).
@@ 5 Proofs of non-existence results
Let $\{Z_{n}\}_{n\in\mathbbm{N}_{0}}$ and $\{X_{n}\}_{n\in\mathbbm{N}_{0}}$ be two Markov chains taking values on $\mathbbm{N}_{0}$. For $j,k\in\mathbbm{N}_{1}$ such that $j<k$, if there exists a monotonically decreasing function $\boldsymbol{\mathfrak{K}}:[M,\infty)\to\mathbbm{R}_{\geq 0}$, $M\in\mathbbm{N}_{0}$, such that
$\mathcal{L}_{(Z_{j+1},\dots,Z_{k})|Z_{j}=u_{j}}-\mathcal{L}_{(X_{j+1},\dots,X_{k})|X_{j}=u_{j}}||_{TV}\leq\boldsymbol{\mathfrak{K}}(u_{j})\quad\forall u_{j}\geq M,$ (16)
then we have
[…]
@@
Proof.
We write $p_{Z_{n}}(i,j):=\mathbbm{P}(Z_{n}=j|Z_{n-1}=i)$ and $p_{X_{n}}(i,j):=\mathbbm{P}(X_{n}=j|X_{n-1}=i)$. From the sum representation of the total variation distance (see [24, Proposition 4.2]) and the triangle inequality,
[…]
@@
[…]
@@
$\sum_{u_{j}=N+1}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\cdot\boldsymbol{\mathfrak{K}}(u_{j})\leq\boldsymbol{\mathfrak{K}}(N+1)\cdot\sum_{u_{j}=N+1}^{\infty}p_{Z_{j}}(u_{j-1},u_{j})\leq\boldsymbol{\mathfrak{K}}(N+1).\\$
Combining these two inequalities then yields our desired result.
@@
for constants $n\in\mathbbm{N}_{1}$, $c_{1},\dots,c_{n}>0$, $q_{1},\dots,q_{n}>0$, and $b>1$. Then $\lim_{z\to\infty}\lim_{k\to\infty}f_{k}(z)=0$.
Proof.
For a given $k\in\mathbbm{N}_{1}$, we can expand $f_{k}$ iteratively to see that
$f_{k}(z)=\sum_{i=1}^{n}\frac{c_{i}}{z^{q_{i}}}\cdot\sum_{j=0}^{k-2}\big{(}b^{-q_{i}}\big{)}^{j}+f_{1}(b^{k-1}\cdot z)=\sum_{i=1}^{n}\frac{c_{i}}{z^{q_{i}}}\cdot\frac{1-b^{-(k-1)q_{i}}}{1-b^{-q_{i}}}+f_{1}(b^{k-1}\cdot z).$
Since $b>1$ and $q_{i}>0$ for all $1\leq i\leq n$, each $b^{-q_{i}}<1$. Therefore, given that $\lim_{x\to\infty}f_{1}(x)=0$, we have that
@@
from which our desired result follows by letting $z\to\infty$.
@@
Proof of Lemma 3.1..
Since $\{Z_{n},z_{0}\}$ is assumed to be supercritical, $\liminf_{z\to\infty}\tau_{Z}(z)>1$. Hence for any $t$ such that $1<t<\liminf_{z\to\infty}\tau_{Z}(z)$, there exists $M_{1}\in\mathbbm{N}_{1}$ such that for all $z\geq M_{1}$, $\varepsilon(z)\cdot m>t\cdot z$. In addition, if $\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})$, then there exists $s>0$ such that for all $z_{0}\geq M_{2}$, $M_{2}\in\mathbbm{N}_{1}$, $\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}\leq s\cdot z_{0}^{-q}$.
@@
$\mathcal{L}_{(Z_{j},\dots,Z_{k})|Z_{j-1}=u_{j-1}}-\mathcal{L}_{(X_{j},\dots,X_{k})|X_{j-1}=u_{j-1}}||_{TV}\leq\boldsymbol{\mathfrak{K}}_{k-j+1}(u_{j-1}),$ (17)
for a decreasing function $\boldsymbol{\mathfrak{K}}_{k-j+1}:[M_{3},\infty)\to\mathbbm{R}_{\geq 0}$ given by $\boldsymbol{\mathfrak{K}}_{1}(z):=s\cdot z^{-q}$ and for $j<k$, $\boldsymbol{\mathfrak{K}}_{k-j+1}(z):=s\cdot z^{-q}+\frac{(a\sigma^{2}+bm^{2})}{(1-\alpha)^{2}t^{2}\cdot z}+\boldsymbol{\mathfrak{K}}_{k-j}(\alpha t\cdot z)$, where $a$ and $b$ are given in (6), and where $\alpha\in(1/t,1)$.
@@
$\mathcal{L}_{Z_{k}|Z_{k-1}=u_{k-1}}-\mathcal{L}_{X_{k}|X_{k-1}=u_{k-1}}||_{TV}\leq\frac{s}{u_{k-1}^{q}}\qquad\textrm{for any $u_{k-1}\geq M_{3}$.}$